Let X 1 , X 2 , ⋯,Xn be a random sample from a normal distribution with mean µ 

 and variance σ2. Consider e X as an estimator of e µ where X  is the sample  

µ mean. Show that e X is consistent estimator of e .




1
Expert's answer
2021-05-27T10:53:44-0400

Showing that X is a consistent ( unbiased ) estimator for u.

"E(estimator )= E(\frac{1}{n}\sum X)"

"=\frac{1}{n}\sum E(x)"

="\frac{1}{n}\sum u"

"\frac{1}{n} n.u"

Therefore "E( estimator )= u"

For clarity on the missing information in the equations see the picture attached.